A beam of white light is incident normally on a plane surface absorbing 70% of the light and reflecting the rest. If the incident beam carries 10 W of power, find the force exerted by it on the surface.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Power 10 W P →Momentum λ = h/c or, h/λ or,P/t = hh/λt R = hc/λ or, E/t = hc/λt = Power (W) W = Pc/t or, P/t = W/c = force. or Force = 7/10 (absorbed) + 2 * 3/10 (reflected) = 7/10 * W/C + 2 * 3/10 * W/C ⇒ 7/10* 10/3 * 10^8 + 2 * 3/10 * 10^3 * 10^8